# High-dimensional Gaussian and bootstrap approximations for robust means

### v3

@@ Appendix E Proofs for Section 4
Consider the first the case of
$C_{1}\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\leq\frac{b_{1}^{2}}{2},$ (E.9)
where $C_{1}$ is the constant from Theorem 3.1 depending only on $b_{2},c$, and $m$. Thus, by that theorem there exists a set $E_{n}$ of probability at least $1-\frac{24}{n}$ on which
$\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}\leq C_{1}\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\leq\frac{b_{1}^{2}}{2}.$
@@
which is well-defined on $E_{n}$. Thus, by the mean-value theorem, there exists a constant $C_{2}$ depending only on $b_{1},b_{2},c$, and $m$ such that on $E_{n}$.
$\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,0,j,k}-\Sigma_{0,j,k}\bigr\rvert}\leq C_{2}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr)}.$
@@
$B_{n}\leq C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\Biggr)}^{1/2}$
for a constant $C$ depending only on $b_{1},b_{2},c$, and $m$ which implies the bound on $\tilde{\rho}_{n,W,S}$ in (15) in case of (E.9).
In case (E.9) is not satisfied, we conclude as in the end of the proof of Theorem 2.1. Finally, $\tilde{\rho}_{n,W,S}\to 0$ by (B.7)–(B.10) of Lemma B.4. ∎
@@ Appendix F Proofs for Section 5
Fix $j\in\mathinner{\{1,\ldots,d\}}$.
Since $\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j})=\tilde{X}_{i,j}$ for $i\in I_{n}=\mathinner{\bigl\{\lceil\varepsilon_{n}n\rceil,\ldots,\lceil(1-\varepsilon_{n})n\rceil\bigr\}}$ one has
@@
$\leq\frac{2\lceil{\varepsilon_{n}n}\rceil}{n}\mathinner{\bigl[|\hat{\alpha}_{j}-\mu_{j}|\vee|\hat{\beta}_{j}-\mu_{j}|\bigr]},$
it suffices to bound $|\hat{\alpha}_{j}-\mu_{j}|\vee|\hat{\beta}_{j}-\mu_{j}|$ from above. To this end, by Lemma G.2 and $\hat{\alpha}_{j}-\mu_{j}\leq\hat{\beta}_{j}-\mu_{j}$ it holds with probability at least $1-\frac{2}{dn}$ that
$|\hat{\alpha}_{j}-\mu_{j}|\vee|\hat{\beta}_{j}-\mu_{j}|\leq|Q_{\varepsilon_{n}-c^{-1}\varepsilon_{n}}(X_{1,j})-\mu_{j}|\vee|Q_{1-\varepsilon_{n}+c^{-1}\varepsilon_{n}}(X_{1,j})-\mu_{j}|\leq\frac{\sigma_{m}}{(\varepsilon_{n}-c^{-1}\varepsilon_{n})^{1/m}},$
where the last estimate is by Lemma G.1. Therefore, there exists a constant $C$ depending only on $b_{2},c$, and $m$ such that with probability at least $1-\frac{2}{dn}$,
$A_{n,j}\leq\frac{4\sqrt{n}\lceil{\varepsilon_{n}n}\rceil\sigma_{m}}{n(\varepsilon_{n}-c^{-1}\varepsilon_{n})^{1/m}}\leq C\sqrt{n}\varepsilon_{n}^{1-\frac{1}{m}}\leq C\sqrt{n}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr)}=\mathrel{\mathop{\ordinarycolon}}\overline{A}_{n},$
@@
$-\mathfrak{A}_{n}-C\overline{A}_{n}\sqrt{\log(d)}-\frac{2}{n}\leq\mathbb{P}\mathinner{\bigl(S_{n,T}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z\in H\bigr)},$ (F.2)
@@
The proof is almost identical to that of Theorem 3.2, but is included for completeness. By the triangle inequality $\tilde{\rho}_{n,T}$ is bounded from above by the sum of
@@
First, $\rho_{n,T}\leq\mathfrak{C}_{n}$, where $\mathfrak{C}_{n}$ is the upper bound on $\rho_{n,T}$ in Theorem 5.1. Next, by the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 of Chernozhuokov et al. (2022) (cf. also Proposition 2 in Chernozhukov et al. (2023a)),
$B_{n}\leq C\mathinner{\Bigl(\log^{2}(d)\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}\Bigr)}^{1/2}\leq C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\Biggr)}^{1/2},$
the last inequality holding with probability at least $1-\frac{24}{n}$ by Theorem 3.1 and $C$ being a constant depending on $b_{1},b_{2},c$ and $m$ only. That $\tilde{\rho}_{n,T}\to 0$ in probability follows from (B.7)–(B.9) of Lemma B.4. ∎
@@ Appendix G Auxiliary lemmas
The following standard lemma is Lemma B.2 from Kock and Preinerstorfer (2025). It bounds the difference between the mean and quantile of a distribution of a random variable $Z$ (which is not necessarily continuous).
@@
Fix $j\in\mathinner{\{1,\ldots,d\}}$, $c\in(1,\sqrt{1.5})$$n\in\mathbb{N}$, and let Assumption 2.1 be satisfied. If $\varepsilon_{n}\in(0,1/2)$ with $\varepsilon_{n}$ as in (5), each of (G.2)–(G.5) below holds with probability at least $1-\frac{1}{dn}$:
$\tilde{X}_{\lceil\varepsilon_{n}n\rceil,j}^{*}$ $\geq$ $Q_{\varepsilon_{n}-c^{-1}\varepsilon_{n}}(X_{1,j});$ (G.2)
$\tilde{X}_{\lceil(1-\varepsilon_{n})n\rceil,j}^{*}$ $\geq$ $Q_{1-\varepsilon_{n}-c^{-1}\varepsilon_{n}}(X_{1,j});$ (G.3)
$\tilde{X}_{\lfloor\varepsilon_{n}n\rfloor+1,j}^{*}$ $\leq$ $Q_{\varepsilon_{n}+c^{-1}\varepsilon_{n}}(X_{1,j});$ (G.4)
$\tilde{X}_{\lfloor(1-\varepsilon_{n})n\rfloor+1,j}^{*}$ $\leq$ $Q_{1-\varepsilon_{n}+c^{-1}\varepsilon_{n}}(X_{1,j}).$ (G.5)
@@
Apply Lemma B.4 in Kock and Preinerstorfer (2025) with $\delta=\frac{6}{dn}$ (recall that $dn>6$ is assumed throughout; cf. the sentence right before Theorem 2.1), $\eta=\overline{\eta}_{n}$ to each coordinate $j=1,\ldots,d$ separately noting that $\varepsilon_{c}$ there equals $\varepsilon_{n}$ for $\delta=\frac{6}{dn}$ and that our Assumption 2.1 implies the assumptions there. ∎
@@
If $\varepsilon_{n}\in(0,1/2)$ with $\varepsilon_{n}$ as in (5), $c\in(1,\sqrt{1.5})$, and Assumption 2.1 is satisfied then, for all $j=1,\ldots,d$,
$\mathinner{\!\biggl\lvert\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(\tilde{X}_{i,j})-\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{i,j})\bigr]}\biggr\rvert}\leq 2\mathinner{\biggl[\frac{1-\sqrt{2(c^{2}-1)}}{(c-1)}\biggr]}^{1/m}\sigma_{m}\overline{\eta}_{n}^{1-\frac{1}{m}}$ (G.6)
@@
$\mathinner{\!\biggl\lvert\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(\tilde{X}_{i,j})-\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})\bigr]}\biggr\rvert}\leq 2\mathinner{\biggl[\frac{1-\sqrt{2(c^{2}-1)}}{(c-1)}\biggr]}^{1/m}\sigma_{m}\overline{\eta}_{n}^{1-\frac{1}{m}}.$ (G.7)
@@
Fix $j\in\mathinner{\{1,\ldots,d\}}$ and recall the definitions of $\underline{\alpha}_{j},\overline{\alpha}_{j},\underline{\beta}_{j}$ and $\overline{\beta}_{j}$ from (A.1) and (A.2). The lemma now follows from applying Lemma B.5 in Kock and Preinerstorfer (2025) to $X_{1,j}$ with $\varepsilon=\varepsilon_{n}$$a=c^{-1}\varepsilon_{n}$, recalling that $\sigma_{m,j}\leq\sigma_{m}$ and using that $\varepsilon_{n}\geq\lambda_{1,c}\overline{\eta}_{n}=\frac{c}{1-\sqrt{2(c^{2}-1)}}\cdot\overline{\eta}_{n}$, noting that our Assumption 2.1 implies the assumptions there
@@
If $\varepsilon_{n}\in(0,1/2)$ with $\varepsilon_{n}$ as in (5), $c\in(1,\sqrt{1.5})$, and Assumption 2.1 is satisfied, then for all $j=1,\ldots,d$,
$\mathinner{\!\bigl\lvert\mathbb{E}\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})-\mu_{j}\bigr\rvert}\leq\sigma_{m}\mathinner{\biggl(2\mathinner{\Bigl(\frac{c-1}{c}\Bigr)}^{1-\frac{1}{m}}+\mathinner{\biggl(1+\mathinner{\Bigl[\frac{c+1}{c-1}\Bigr]}^{\frac{1}{m}}\biggr)}\mathinner{\Bigl(\frac{c+1}{c}\Bigr)}^{1-\frac{1}{m}}\biggr)}\varepsilon_{n}^{1-\frac{1}{m}}$ (G.8)
@@
$\mathinner{\!\bigl\lvert\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1})-\mu_{j}\bigr\rvert}\leq\sigma_{m}\mathinner{\biggl(2\mathinner{\Bigl(\frac{c-1}{c}\Bigr)}^{1-\frac{1}{m}}+\mathinner{\biggl(1+\mathinner{\Bigl[\frac{c+1}{c-1}\Bigr]}^{\frac{1}{m}}\biggr)}\mathinner{\Bigl(\frac{c+1}{c}\Bigr)}^{1-\frac{1}{m}}\biggr)}\varepsilon_{n}^{1-\frac{1}{m}}.$ (G.9)

### v4

@@ Appendix E Proofs for Section 4
Arguing as in the opening paragraph of the proof of Theorem 4.1, it suffices to consider the case of (E.3) being satisfied.
Let $C_{1}$ be the constant from Theorem 3.1 depending only on $b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$. By that theorem (note that (9) is ensured to be true by (E.3) so that Theorem 3.1 can be employed) there exists a set $E_{n}$ of probability at least $1-\frac{24}{n}$ on which
$\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}\leq C_{1}\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\leq\frac{b_{1}^{2}}{2}.$
@@
which is well-defined on $E_{n}$. Thus, by the mean-value theorem, there exists a constant $C_{2}$ depending only on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$ such that on $E_{n}$.
$\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,0,j,k}-\Sigma_{0,j,k}\bigr\rvert}\leq C_{2}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}.$
@@
$B_{n}\leq C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}^{1/2}$
for a constant $C$ depending only on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$ which implies the bound on $\tilde{\rho}_{n,W,S}$ in (13) in case of (E.3).
Finally, $\tilde{\rho}_{n,W,S}\to 0$ by (B.7)–(B.10) of Lemma B.4. ∎
@@ Appendix F Proofs for Section 5
Since $\log(dn)/n\leq\lambda_{2}^{-1}\varepsilon_{n}$, one has that $\mathfrak{T}_{n}^{(1)}$ in (9) is bounded as
$\mathfrak{T}_{n}^{(1)}\leq 2\varepsilon_{n}+\lambda_{2}^{-1}\varepsilon_{n}+\sqrt{\lambda_{2}^{-2}\varepsilon_{n}^{2}+4\lambda_{2}^{-1}\varepsilon_{n}^{2}}\leq C_{1}\varepsilon_{n},$
for a constant $C_{1}$ depending only on $\lambda_{2}$. In the sequel we impose that $C_{1}\varepsilon_{n}<1/2$ such that by the previous display (A.5) is satisfied allowing us to invoke Lemma G.3 below. In case $C_{1}\varepsilon_{n}\geq 1/2$, (15) trivially holds by the same arguments as in the end of the proof of Theorem 2.1.
Fix $j\in\mathinner{\{1,\ldots,d\}}$. Since $\phi_{\hat{\alpha}_{j},\hat{\beta}_{j}}(\tilde{X}_{i,j})=\tilde{X}_{i,j}$ for $i\in I_{n}=\mathinner{\bigl\{\lceil\varepsilon_{n}n\rceil,\ldots,\lceil(1-\varepsilon_{n})n\rceil\bigr\}}$
@@
$=\frac{2\lceil{\varepsilon_{n}n}\rceil}{n}\mathinner{\bigl[|\hat{\alpha}_{j}-\mu_{j}|\vee|\hat{\beta}_{j}-\mu_{j}|\bigr]},$
it suffices to bound $|\hat{\alpha}_{j}-\mu_{j}|\vee|\hat{\beta}_{j}-\mu_{j}|$ from above. To this end, by Lemma G.3 and $\hat{\alpha}_{j}-\mu_{j}\leq\hat{\beta}_{j}-\mu_{j}$ ((A.5) implies $\varepsilon_{n}<1/4$ such that $\hat{\alpha}_{j}=\tilde{X}_{\lceil{\varepsilon_{n}n}\rceil,j}^{*}\leq\tilde{X}_{\lceil(1-\varepsilon_{n})n\rceil,j}^{*}=\hat{\beta}_{j}$) it holds with probability at least $1-\frac{2}{dn}$ that
$|\hat{\alpha}_{j}-\mu_{j}|\vee|\hat{\beta}_{j}-\mu_{j}|\leq|Q_{c_{1,n}\varepsilon_{n}}(X_{1,j})-\mu_{j}|\vee|Q_{1-c_{1,n}\varepsilon_{n}}(X_{1,j})-\mu_{j}|\leq\frac{\sigma_{m}}{(c_{1,n}\varepsilon_{n})^{1/m}},$
where the last estimate is by Lemma G.1. Therefore, by Lemma G.2 there exists a constant $C$ depending only on $b_{2},\lambda_{1},\lambda_{2}$, and $m$ such that with probability at least $1-\frac{2}{dn}$,
$A_{n,j}\leq\frac{4\sqrt{n}\lceil{\varepsilon_{n}n}\rceil\sigma_{m}}{n(c_{1,n}\varepsilon_{n})^{1/m}}\leq C\sqrt{n}\varepsilon_{n}^{1-\frac{1}{m}}\leq C\sqrt{n}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr)}=\mathrel{\mathop{\ordinarycolon}}\overline{A}_{n},$
@@
$\mathfrak{A}_{n}-C\overline{A}_{n}\sqrt{\log(d)}-\frac{2}{n}\leq\mathbb{P}\mathinner{\bigl(S_{n,T}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(Z\in H\bigr)},$ (F.2)
@@
The proof is almost identical to that of Theorem 3.2, but is included for completeness. Note first that writing $\mathfrak{T}_{n}^{(3)}$ for the left-hand side of (9), it follows from $\frac{\log(d^{2}n)}{n}\leq{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}$ that
$\mathfrak{T}_{n}^{(3)}\leq 2\varepsilon_{n}^{\prime}+{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}+\sqrt{{\lambda_{2}^{\prime}}^{-2}{\varepsilon_{n}^{\prime}}^{2}+4{\lambda_{2}^{\prime}}^{-1}{\varepsilon_{n}^{\prime}}^{2}}\leq C_{1}\varepsilon_{n}^{\prime},$
for a constant $C_{1}$ depending only on $\lambda_{2}^{\prime}$.
Suppose first that $C_{1}\varepsilon_{n}^{\prime}<1$ such that (9) is satisfied by the previous display. By the triangle inequality $\tilde{\rho}_{n,T}$ is bounded from above by the sum of
@@
First, $\rho_{n,T}\leq\mathfrak{C}_{n}$, where $\mathfrak{C}_{n}$ is the upper bound on $\rho_{n,T}$ in Theorem 5.1. Next, by the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 of Chernozhukov et al. (2022) (cf. also Proposition 2 in Chernozhukov et al. (2023a)),
$B_{n}\leq C\mathinner{\Bigl(\log^{2}(d)\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\tilde{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}\Bigr)}^{1/2}\leq C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{1}{(m/2)\wedge 2}}\biggr]}\Biggr)}^{1/2},$
the last inequality holding with probability at least $1-\frac{24}{n}$ by Theorem 3.1 and $C$ being a constant depending on $b_{1},b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$ and $m$ only.
If, on the other hand, $C_{1}\varepsilon_{n}^{\prime}\geq 1$ then by the definition of $\varepsilon_{n}^{\prime}$ in (7)
$\overline{\eta}_{n}\geq\frac{1}{2C_{1}\lambda_{1}^{\prime}}\qquad\text{or}\qquad\frac{\log(dn)}{n}=\frac{\log([dn]^{2})/2}{n}\geq\frac{\log(d^{2}n)}{2n}\geq\frac{1}{4C_{1}\lambda_{2}^{\prime}}.$
In either case, the upper bound on $\tilde{\rho}_{n,T}$ remains valid by suitably adjusting $C$ (if needed).
That $\tilde{\rho}_{n,T}\to 0$ in probability follows from (B.7)–(B.9) of Lemma B.4. ∎
@@ Appendix G Auxiliary lemmas
The following standard result is Lemma C.1 from Kock and Preinerstorfer (2025). It bounds the difference between the mean and quantile of a distribution of a random variable $Z$ (which is not necessarily continuous).
@@
Lemma G.2.
For $c_{1,n}$ as defined in (A.1), $c_{2,n}$ as defined in (A.2), and $\varepsilon_{n}$ as defined in (4) and satisfying (A.5), it holds that
$0<(1-\lambda_{1}^{-1})\exp\mathinner{\Bigl({-\frac{1}{\lambda_{2}(1-\lambda_{1}^{-1})}-1}\Bigr)}\leq\$$c_{1,n}<1,$
$1<\$$c_{2,n}\leq 2+\lambda_{2}^{-1}+\sqrt{\lambda_{2}^{-2}+4\lambda_{2}^{-1}},$
and
$0<\varepsilon_{n}\min(c_{1,n},c_{2,n})$ $\leq\varepsilon_{n}(c_{1,n}+c_{2,n})$
$\leq 2\varepsilon_{n}+\frac{\log(dn)}{n}+\sqrt{\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{2}+4\frac{\log(dn)}{n}\varepsilon_{n}}.$ (G.2)
Proof.
Apply Lemma B.3 in Kock and Preinerstorfer (2025) with $\delta=\frac{6}{dn}$ (recall that $dn>6$ is assumed throughout; cf. the sentence right before Theorem 2.1), $\epsilon=\varepsilon_{n}$, and $\eta=\overline{\eta}_{n}$. To this end, note that $c_{1}$, and $c_{2}$ there equal $c_{1,n}$ and $c_{2,n}$, respectively, for $\delta=\frac{6}{dn}$, that $\varepsilon_{n}\geq\lambda_{2}\log(dn)/n$ by definition, and that $\varepsilon_{n}<1/4$ follows from (A.5). ∎
@@
Fix $j\in\mathinner{\{1,\ldots,d\}}$$n\in\mathbb{N}$, let Assumption 2.1 be satisfied, and $\varepsilon_{n}$ be as defined in (4) satisfying (A.5). Then, for $c_{1,n}$ as in (A.1) and $c_{2,n}$ as in (A.2), we have
$0<\varepsilon_{n}\min(c_{1,n},c_{2,n})\leq\varepsilon_{n}(c_{1,n}+c_{2,n})<1/2,$
and each of (G.3)–(G.6) below holds with probability at least $1-\frac{1}{dn}$:
$\tilde{X}_{\lceil\varepsilon_{n}n\rceil,j}^{*}$ $\geq$ $Q_{c_{1,n}\varepsilon_{n}}(X_{1,j});$ (G.3)
$\tilde{X}_{\lceil(1-\varepsilon_{n})n\rceil,j}^{*}$ $\geq$ $Q_{1-c_{2,n}\varepsilon_{n}}(X_{1,j});$ (G.4)
$\tilde{X}_{\lfloor\varepsilon_{n}n\rfloor+1,j}^{*}$ $\leq$ $Q_{c_{2,n}\varepsilon_{n}}(X_{1,j});$ (G.5)
$\tilde{X}_{\lfloor(1-\varepsilon_{n})n\rfloor+1,j}^{*}$ $\leq$ $Q_{1-c_{1,n}\varepsilon_{n}}(X_{1,j}).$ (G.6)
@@
The first statement in the present lemma follows from (A.5) along with (G.2) of Lemma G.2. For the remaining statements, apply Lemma B.5 in Kock and Preinerstorfer (2025) with $\delta=\frac{6}{dn}$ (recall that $dn>6$ is assumed throughout; cf. the sentence right before Theorem 2.1), $\epsilon=\varepsilon_{n}$, and $\eta=\overline{\eta}_{n}$. To this end, note that
  1. 1.
    $c_{1}$, and $c_{2}$ there equal $c_{1,n}$ and $c_{2,n}$, respectively, for $\delta=\frac{6}{dn}$,
  2. 2.
    that $\varepsilon_{n}\geq\lambda_{1}\overline{\eta}_{n}$ by definition and $\varepsilon_{n}c_{2,n}<1/2$ (which has already been verified),
  3. 3.
    $\varepsilon_{n}<1/4$ follows from (A.5),
and that our Assumption 2.1 implies the remaining assumptions there. ∎
@@
There exists a positive constant $C$ depending only on $\lambda_{1},\lambda_{2}$, and $m$, such that if $\varepsilon_{n}$ is as in (4) satisfying (A.5) and Assumption 2.1 is satisfied, then for all $j=1,\ldots,d$,
$\mathinner{\!\biggl\lvert\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(\tilde{X}_{i,j})-\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{i,j})\bigr]}\biggr\rvert}\leq C\sigma_{m}\overline{\eta}_{n}^{1-\frac{1}{m}}$ (G.7)
@@
$\mathinner{\!\biggl\lvert\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(\tilde{X}_{i,j})-\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{i,j})\bigr]}\biggr\rvert}\leq C\sigma_{m}\overline{\eta}_{n}^{1-\frac{1}{m}}.$ (G.8)
@@
Fix $j\in\mathinner{\{1,\ldots,d\}}$ and recall the definitions of $\underline{\alpha}_{j},\overline{\alpha}_{j},\underline{\beta}_{j}$ and $\overline{\beta}_{j}$ from (A.3) and (A.4). To establish (G.7), apply Lemma C.2 in Kock and Preinerstorfer (2025) to $X_{1,j}$ with $\eta=\overline{\eta}_{n}$$s_{1}=c_{1,n}\varepsilon_{n}$, and $s_{2}=1-c_{2,n}\varepsilon_{n}$, recalling that $\sigma_{m,j}\leq\sigma_{m}$ and using that $\varepsilon_{n}>\lambda_{1}\overline{\eta}_{n}>\overline{\eta}_{n}$ as well as the lower bounds on $c_{1,n}$ and $c_{2,n}$ from Lemma G.2, noting that our Assumption 2.1 implies the assumptions there. The requirement $s_{1}<s_{2}$ in Lemma C.2 in Kock and Preinerstorfer (2025) follows by (A.5) and Lemma G.2. (G.8) follows in a similar fashion using $s_{1}=c_{2,n}\varepsilon_{n}$ and $s_{2}=1-c_{1,n}\varepsilon_{n}$ instead. ∎
@@
There exists a positive constant $C$ depending only on $\lambda_{1},\lambda_{2}$, and $m$, such that if $\varepsilon_{n}$ is as in (4) satisfying (A.5) and Assumption 2.1 is satisfied, then for all $j=1,\ldots,d$,
$\mathinner{\!\bigl\lvert\mathbb{E}\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})-\mu_{j}\bigr\rvert}\vee\mathinner{\!\bigl\lvert\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1})-\mu_{j}\bigr\rvert}\leq C\sigma_{m}\varepsilon_{n}^{1-\frac{1}{m}}$ (G.9)
Proof.
Fix $j\in\mathinner{\{1,\ldots,d\}}$ and recall the definitions of $\underline{\alpha}_{j},\overline{\alpha}_{j},\underline{\beta}_{j}$ and $\overline{\beta}_{j}$ from (A.3) and (A.4). Since $\mathbb{E}\phi_{\underline{\alpha}_{j},\underline{\beta}_{j}}(X_{1,j})-\mu_{j}\leq\mathbb{E}\phi_{\overline{\alpha}_{j},\overline{\beta}_{j}}(X_{1,j})-\mu_{j}$, (G.9) follows from i) (C.5) of Lemma C.5 in Kock and Preinerstorfer (2025) applied to $X_{1,j}$ with $s_{1}=c_{1,n}\varepsilon_{n}$$s_{2}=1-c_{2,n}\varepsilon_{n}$ and ii) (C.6) of Lemma C.5 in Kock and Preinerstorfer (2025) applied to $X_{1,j}$ with $s_{1}=c_{2,n}\varepsilon_{n}$$s_{2}=1-c_{1,n}\varepsilon_{n}$, recalling that $\sigma_{m,j}\leq\sigma_{m}$ (noting that our Assumption 2.1 implies the assumptions there, and in either case the requirement $s_{1}<s_{2}$ imposed in Lemma C.5 in Kock and Preinerstorfer (2025) follows from (A.5) and Lemma G.2), such that
$\leq 2\sigma_{m}(c_{1,n}\varepsilon_{n})^{1-\frac{1}{m}}+\sigma_{m}\mathinner{\Bigl(1+\mathinner{\Bigl[\frac{c_{2,n}\varepsilon_{n}}{1-c_{2,n}\varepsilon_{n}}\Bigr]}^{\frac{1}{m}}\Bigr)}(c_{2,n}\varepsilon_{n})^{1-\frac{1}{m}}$
$\leq 2\sigma_{m}(c_{1,n}\varepsilon_{n})^{1-\frac{1}{m}}+2\sigma_{m}(c_{2,n}\varepsilon_{n})^{1-\frac{1}{m}},$
the second inequality following from $c_{2,n}\varepsilon_{n}<1/2$ by Lemma G.2 and (A.5) and $(0,1)\ni z\mapsto z/(1-z)$ being increasing. The upper bounds on $c_{1,n}$ and $c_{2,n}$ from Lemma G.2 now yield the conclusion of the lemma. ∎
The following result is similar to Lemma G.2 and bounds $c_{1,n}^{\prime}$ and $c_{2,n}^{\prime}$ as defined in (D.1) and (D.2), respectively. It will be used in Lemma (G.7) below, as well as in the proof of Theorem 3.1.
Lemma G.6.
For $c_{1,n}^{\prime}$ as defined in (D.1), $c_{2,n}^{\prime}$ as defined in (D.2), and $\varepsilon_{n}^{\prime}$ as defined in (7) and satisfying (9), it holds that
$0<(1-\lambda_{1}^{\prime-1})\exp\mathinner{\Bigl({-\frac{1}{\lambda_{2}^{\prime}(1-\lambda_{1}^{\prime-1})}-1}\Bigr)}\leq\$$c_{1,n}^{\prime}<1,$
$1<\$$c_{2,n}^{\prime}\leq 2+\lambda_{2}^{\prime-1}+\sqrt{\lambda_{2}^{\prime-2}+4\lambda_{2}^{\prime-1}},$
@@
$0<\varepsilon_{n}^{\prime}\min(c_{1,n}^{\prime},c_{2,n}^{\prime})$ $\leq\varepsilon_{n}^{\prime}(c_{1,n}^{\prime}+c_{2,n}^{\prime})$
$\leq 2\varepsilon_{n}^{\prime}+\frac{\log(d^{2}n)}{n}+\sqrt{\mathinner{\Bigl(\frac{\log(d^{2}n)}{n}\Bigr)}^{2}+4\frac{\log(d^{2}n)}{n}\varepsilon_{n}^{\prime}}.$ (G.10)